The American Baker’s Association reports that annual sales of bakery goods last year rose 15 percent, driven by a 50 percent increase in the demand for bran muffins. Most of the increase was attributed to a report that diets rich in bran help prevent certain types of cancer. You are the manager of a bakery that produces and packages gourmet bran muffins, and you currently sell bran muffins in packages of three. However, as a result of this new report, a typical consumer’s inverse demand for your bran muffins is now P = 10 - 2Q. If your cost of producing bran muffins is C(Q) = 2.0Q, determine the optimal number of bran muffins to sell in a single package and the optimal package price.

Instruction: Enter your response for the optimal package price rounded to two decimal places.

Optimal package size: units

Optimal package price: $
Answer #1

Solution:

We are given the new demand function as: P = 10 - 2Q

Then, total revenue = price*quantity

TR = (10 - 2Q)*Q

TR = 10Q - 2Q2

Then, marginal revenue, MR = dTR/dQ

MR = 10 - 2*2Q so, MR = 10 - 4Q

Cost function is given as C(Q) = 2.0*Q

So, marginal cost = dC(Q)/dQ = 2

Optimal quantity occurs where MR = MC

So, 10 - 4Q = 2

Q = (10 - 2)/4 = 2

And, corresponding price is P = 10 - 2*2 = $6

So, optimal package is 2 units (or now sales would be bran muffin package of two), and optimal price per package is $6.